# Beyond Superficial Forgetting: Thorough Unlearning through Knowledge Density Estimation and Block Re-insertion

### v1

@@ Abstract
Machine unlearning, which selectively removes harmful knowledge from a pre-trained model without retraining from scratch, is crucial for addressing privacy, regulatory compliance, and ethical concerns in Large Language Models (LLMs). However, existing unlearning methods often struggle to thoroughly remove harmful knowledge, leaving residual harmful knowledge that can be easily recovered. To address these limitations, we propose Knowledge Density-Guided Unlearning via Blocks Reinsertion (KUnBR), a novel approach that first identifies layers with rich harmful knowledge and then thoroughly eliminates the harmful knowledge via re-insertion strategy. Our method introduces knowledge density estimation to quantify and locate layers containing the most harmful knowledge, enabling precise unlearning. Additionally, we design a layer re-insertion strategy that extracts and re-inserts harmful knowledge-rich layers into the original LLM, bypassing gradient obstruction caused by cover layers and ensuring effective gradient propagation during unlearning. Extensive experiments conducted on several unlearning and general capability benchmarks demonstrate that KUnBR achieves state-of-the-art forgetting performance while maintaining model utility.
@@ Introduction
Machine unlearning (liu2025rethinking; 9519428) refers to selectively removing specific subsets of knowledge, such as privacy-sensitive or harmful content, from a pre-trained model without retraining it from scratch (carlini2021extractingtrainingdatalarge; xu2024machine). This task has become increasingly crucial for the development of large language models (LLMs) (gpt4o; llama3; claude35sonnet; guo2025deepseek), as it addresses growing concerns around data privacy (carlini2021extractingtrainingdatalarge; huang2022largepretrainedlanguagemodels; lee2024protecting; liu2024learning) and the ethical issue of AI systems (2021On). Unlearning is critical not only for addressing regulatory requirements such as “right to be forgotten”, but also for ensuring that LLMs remain secure, reliable, and aligned with societal values.
@@
Although existing machine unlearning methods can suppress harmful knowledge, several jailbreak attack studies  (zhou2024don; liu2023prompt; schwinn2024soft; rimsky2024steering) have shown that the robustness issues remain. The Retraining on T (RTT) (deeb2025unlearningmethodsremoveinformation), which is an attack method at the parameter-modification level, demonstrates that minimal re-training on a subset (a.k.a., the T set) of the forget set can restore most of the supposedly eliminated knowledge. These results demonstrate that the model parameters still contain a substantial amount of knowledge that should have been forgotten, which reveals the inability of existing methods to thoroughly remove knowledge from the model parameters. That means that existing methods often rely on the adjustment of a small number of model parameters (a.k.a., cover layers) to mask or suppress the representation of harmful knowledge, merely preventing the model from outputting undesired content without truly eliminating it from the model’s internal representations. This fundamental limitation suggests the need for more robust and thorough unlearning methods.
In this paper, we propose Knowledge Density-Guided Unlearning via Blocks Reinsertion (KUnBR), which identifies blocks with rich harmful knowledge, and iteratively performs independent unlearning on these blocks via re-insertion strategy, enables a deeper level of unlearning. We first introduce a knowledge density estimation method to identify the layers that contain the most harmful knowledge. By calculating the absolute value of gradients associated with the forget set, knowledge density estimation can locate layers containing high-density knowledge. Instead of superficially modifying cover layer parameters to suppress model outputs, we propose a novel re-insertion strategy to thoroughly unlearn undesired knowledge. This approach extracts knowledge-rich blocks (selected according to the knowledge density estimation) from the unlearned LLM and re-inserts them into the original LLM without conducting the unlearning training. We then apply the unlearning method to train this “grafted” model, which contains the re-inserted layers, with a focus on deeper removal of the undesired knowledge left due to the influence of cover layers. By bypassing the obstruction of cover layers, this strategy ensures more effective gradient propagation and enhances the model’s ability to forget. This simple but efficient strategy significantly reduces the vulnerability of the model to attacks like RTT, which exploit the residual knowledge left by conventional unlearning methods. Extensive experiments conducted on WMDP-Deduped, Years, Random Birthdays and RKWU benchmark datasets demonstrate that our method achieves state-of-the-art performance, since it can remove harmful knowledge more thoroughly and suppress knowledge recovery caused by RTT attack methods.
@@
Figure 2 shows the architecture of KUnBR. The first step is the global “warm-up” unlearning phase, in which we apply a standard Gradient Difference method to adjust all model parameters at once; In the second step, we perform knowledge density estimation and our block-selection strategy to pick out those blocks that contain high-density knowledge. Finally, we introduce a re-insertion strategy to bypass the masking effect of cover layers and enable any remaining knowledge to be further eliminated.
@@
In this paper, we propose a block selection strategy to dynamically select blocks for unlearning. Table 4 presents the performance on the Years dataset when using different numbers of blocks and Top-$K$. We observe that using fewer blocks (e.g., M=4M{=}4) hampers forgetting, as each block spans many layers, making it difficult to isolate and erase specific knowledge. This often leads to premature convergence and weaker RTT resistance. Conversely, using too many blocks (e.g., M=32M{=}32) may overlook inter-layer dependencies that are essential for effective unlearning. Overall, Table 4 shows that most $(M,K)$ configurations substantially outperform the baselines. Remarkably, the same configuration performs well on both LLaMA-3 and Zephyr, suggesting strong cross-architecture stability. This indicates that users can safely adopt the default configuration without additional re-tuning.
We also acknowledge a potential concern that shallow forgetting might simply shift to a finer granularity. However, we interpret layer blocks as constituent memory units rather than isolated mini-language models, consistent with the finding that Transformer Feed-Forward Layers Are Key-Value Memories. Furthermore, even when we reduce the block granularity down to the layer level—the smallest structural unit of the model—our method still maintains strong unlearning robustness. We attribute this to genuine knowledge removal rather than mere inaccessibility, confirming that our approach achieves true forgetting without relying on superficial compression effects.
@@
To fairly compare the training efficiency of different methods, we report the time cost of the unlearning phase on the Years dataset, excluding evaluation overhead. All experiments are conducted on a single NVIDIA A800 GPU. The training times (in minutes) are: GA (24), GD (20), RMU (9), RIA (8), NPO (16), and our method KUnBR (17). From these results, it can be observed that the time consumption of our proposed KUnBR is comparable to other baseline methods. Although it is 88.9% higher than RMU, it is 15% lower than the current state-of-the-art GD method, which further demonstrates the effectiveness of our proposed KUnBR. The reason for this is that, compared to RMU, our method employs block-wise unlearning approach, enabling the model to more thoroughly unlearn harmful knowledge, although at the cost of increased computational complexity. In contrast, the current SOTA baseline GD, due to the formation of cover layers, is unable to achieve complete forgetting of harmful knowledge even with longer training times than KUnBR.

### v2

@@ Abstract
Machine unlearning, which selectively removes harmful knowledge from a pre-trained model without retraining from scratch, is crucial for addressing privacy, regulatory compliance, and ethical concerns in Large Language Models (LLMs). However, existing unlearning methods often struggle to thoroughly remove harmful knowledge, leaving residual harmful knowledge that can be easily recovered. To address these limitations, we propose Knowledge Density-Guided Unlearning via Blocks Reinsertion (KUnBR), a novel approach that first identifies layers with rich harmful knowledge and then thoroughly eliminates the harmful knowledge via re-insertion strategy. Our method introduces knowledge density estimation to quantify and locate layers containing the most harmful knowledge, enabling precise unlearning. Additionally, we design a layer re-insertion strategy that extracts and re-inserts harmful knowledge-rich layers into the original LLM, bypassing gradient obstruction caused by cover layers and ensuring effective gradient propagation during unlearning. Extensive experiments conducted on several unlearning and general capability benchmarks demonstrate that KUnBR achieves state-of-the-art forgetting performance while maintaining model utility [^1] Code available at github.com/llmgfffffff/Beyond-Superficial-Forgetting-KUnBR.
@@ Introduction
Machine unlearning (liu2025rethinking; 9519428) refers to selectively removing specific subsets of knowledge, such as privacy-sensitive or harmful content, from a pre-trained model without retraining it from scratch (carlini2021extractingtrainingdatalarge; xu2024machine). This task has become increasingly crucial for the development of large language models (LLMs) (gpt4o; llama3; claude35sonnet; 10.1145/3589334.3645448; gao2024simulatingfinancialmarketlarge), as it addresses growing concerns around data privacy (carlini2021extractingtrainingdatalarge; huang2022largepretrainedlanguagemodels; lee2024protecting; liu2024learning; 10.1145/3589334.3645448) and the ethical issue of AI systems (2021On). Unlearning is critical not only for addressing regulatory requirements such as “right to be forgotten”, but also for ensuring that LLMs remain secure, reliable, and aligned with societal values.
@@
Although existing machine unlearning methods can suppress harmful knowledge, several jailbreak attack studies  (zhou2024don; liu2023prompt; schwinn2024soft; rimsky2024steering) have shown that the robustness issues remain.
The Retraining on T (RTT) (deeb2025unlearningmethodsremoveinformation), which is an attack method at the parameter-modification level, demonstrates that minimal re-training on a subset (a.k.a., the T set) of the forget set can restore most of the supposedly eliminated knowledge.
These results demonstrate that the model parameters still contain a substantial amount of knowledge that should have been forgotten, which reveals the inability of existing methods to thoroughly remove knowledge from the model parameters. That means that existing methods often rely on the adjustment of a small number of model parameters (a.k.a., cover layers) to mask or suppress the representation of harmful knowledge, merely preventing the model from outputting undesired content without truly eliminating it from the model’s internal representations. This fundamental limitation suggests the need for more robust and thorough unlearning methods.
In this paper, we propose Knowledge Density-Guided Unlearning via Blocks Reinsertion (KUnBR), which identifies blocks with rich harmful knowledge, and iteratively performs independent unlearning on these blocks via re-insertion strategy, enables a deeper level of unlearning.
We first introduce a knowledge density estimation method to identify the layers that contain the most harmful knowledge. By calculating the absolute value of gradients associated with the forget set, knowledge density estimation can locate layers containing high-density knowledge. Instead of superficially modifying cover layer parameters to suppress model outputs, we propose a novel re-insertion strategy to thoroughly unlearn undesired knowledge. This approach extracts knowledge-rich blocks (selected according to the knowledge density estimation) from the unlearned LLM and re-inserts them into the original LLM without conducting the unlearning training. We then apply the unlearning method to train this “grafted” model, which contains the re-inserted layers, with a focus on deeper removal of the undesired knowledge left due to the influence of cover layers. By bypassing the obstruction of cover layers, this strategy ensures more effective gradient propagation and enhances the model’s ability to forget. This simple but efficient strategy significantly reduces the vulnerability of the model to attacks like RTT, which exploit the residual knowledge left by conventional unlearning methods. Extensive experiments conducted on WMDP-Deduped, Years, Random Birthdays and RKWU benchmark datasets demonstrate that our method achieves state-of-the-art performance, since it can remove harmful knowledge more thoroughly and suppress knowledge recovery caused by RTT attack methods.
@@
Figure 2 shows the architecture of KUnBR. The first step is the global “warm-up” unlearning phase, in which we apply a standard Gradient Difference method to adjust all model parameters at once;
In the second step, we perform knowledge density estimation and our block-selection strategy to pick out those blocks that contain high-density knowledge. Finally, we introduce a re-insertion strategy to bypass the masking effect of cover layers and enable any remaining knowledge to be further eliminated.
@@
We propose a block selection strategy to dynamically select blocks for unlearning. Table 4 shows performance on the Years dataset with varying $(M,K)$. Fewer blocks (e.g., M=4M{=}4) hamper forgetting as they span many layers, hindering knowledge isolation and causing premature convergence. Conversely, excessive blocks (e.g., M=32M{=}32) may overlook essential inter-layer dependencies. Most configurations substantially outperform baselines. Notably, the consistency across LLaMA-3 and Zephyr suggests strong cross-architecture stability, allowing users to adopt default settings without re-tuning.
Addressing concerns that shallow forgetting shifts to finer granularity, we interpret blocks as constituent memory units (aligned with FFNs as Key-Value Memories) rather than isolated mini-LMs. Even at the layer level—the smallest unit—our method maintains strong robustness. This confirms genuine knowledge removal rather than mere inaccessibility or superficial compression.
@@
To fairly compare the training efficiency of different methods, we report the time cost of the unlearning phase on the Years dataset, excluding evaluation overhead. All experiments are conducted on a single NVIDIA A800 GPU. The training times (in minutes) are: GA (24), GD (20), RMU (9), RIA (8), NPO (16), and our method KUnBR (17).
From these results, it can be observed that the time consumption of our proposed KUnBR is comparable to other baseline methods. Although it is 88.9% higher than RMU, it is 15% lower than the current state-of-the-art GD method, which further demonstrates the effectiveness of our proposed KUnBR. The reason for this is that, compared to RMU, our method employs block-wise unlearning approach, enabling the model to more thoroughly unlearn harmful knowledge, although at the cost of increased computational complexity. In contrast, the current SOTA baseline GD, due to the formation of cover layers, is unable to achieve complete forgetting of harmful knowledge even with longer training times than KUnBR.